OpenStudy (anonymous):

Latisha earns $5 an hour painting her neighbor's fence. Part A: Create two variables and determine which is dependent and which is independent for this situation. (4 points) Part B: List three ordered pairs to show the values of the two variables. (3 points) Part C: Write an equation to represent the relationship between the two variables. (3 points)

OpenStudy (anonymous):

I have a and c done.

OpenStudy (anonymous):

What is your equation for A? So I can help you with B

OpenStudy (anonymous):

my equation for a is p=5h

OpenStudy (anonymous):

@pflorido are you still helping?

OpenStudy (danjs):

You still working on this one?

OpenStudy (anonymous):

Yes i need help on B

OpenStudy (danjs):

ok

OpenStudy (danjs):

For A) Earned = 5*(Hours)

OpenStudy (danjs):

the independent variable is hours, it determines how much is earned, earned is 'dependent' on hours

OpenStudy (anonymous):

Pay is the dependent and Hours is the independent. p=5h

OpenStudy (anonymous):

i have that already ^

OpenStudy (danjs):

oh i didnt see the dependent/indep

OpenStudy (anonymous):

sorry lol i forgot to post it

OpenStudy (danjs):

Part B, they just want 3 data points from the equation... p = 5h The points are (h , p) pick any values for h you want, 1,2,3....

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

pick one number or 123 is all i need?

OpenStudy (danjs):

any number of hours, and calclulate the earned amount

OpenStudy (anonymous):

what do you mean?

OpenStudy (danjs):

They want 3 points on the graph of y = 5*x Chose any 3 values of X, and determine the corresponding y's.. p = 5*h Let h = 1, then p = 5*1 = 5 Point - (h , p) = (1 , 5)

OpenStudy (danjs):

Let h = 2, then p = 5 * 2 = 10 point (h , p) = (2 , 10)

OpenStudy (danjs):

need one more point

OpenStudy (anonymous):

H=3 p=5*3=15 point (h,p) = (3,15)

OpenStudy (danjs):

right, the last part we did in part A, the equation p = 5h

OpenStudy (anonymous):

so thats all for part b?